OpenStudy (anonymous):

Find the volume of a cone with a base area of 16 pi squared feet and a height of 9 ft. ???

OpenStudy (wolf1728):

When you say the base area is 16 pi squared is that 16 * PI * PI? ( or 157.9136704174)

OpenStudy (anonymous):

Yes

OpenStudy (wolf1728):

Just curious because area of a circel = PI * r² and it seems weird getting PI squared as an answer.

OpenStudy (wolf1728):

Maybe it is 16 PI square feet. That makes it easier to understand.

OpenStudy (anonymous):

haha yes sorry, I messed up

OpenStudy (wolf1728):

area = PI * r² 16 PI = PI * r² r = 4

OpenStudy (anonymous):

how did you figure that out?

OpenStudy (wolf1728):

The area is 16 PI which has to equal PI r² Dividing both sides by PI makes it 16 = r² r = sq root (16) which is 4

OpenStudy (anonymous):

Ok, I got it know Thanks!!!

OpenStudy (wolf1728):

Now for the volume, it equals (PI * r² * h) / 3 =(PI * 16 * 9) / 3 =452.39 / 3 = 150.80 cubic feet

OpenStudy (anonymous):

thanks!(: